OpenStudy (anonymous):

The equation f(x) = 3x2 − 24x + 8 represents a parabola. What is the vertex of the parabola? (−4, 152) (4, −40) (−3, 107) (3, −37)

OpenStudy (anonymous):

will fan and medal!?!??!

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\) in your case \(a=3\) and \(b=-24\) what do you get ?

OpenStudy (anonymous):

i really dont get like any of it

OpenStudy (misty1212):

what i meant was "what do you get for \[-\frac{b}{2a}\]when you replace \(a\) by \(3\) and \(b\) by \(-24\) and do the arithmetic

OpenStudy (anonymous):

3/48

OpenStudy (anonymous):

negative

OpenStudy (misty1212):

not very close actually the b goes up top, the a goes in the bottom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

24/6

OpenStudy (misty1212):

\[\huge -\frac{\color{blue}b}{2\times \color{red}a}\]yeah that looks good

OpenStudy (anonymous):

24/6

OpenStudy (misty1212):

of course 6 goes in to 24 evenly so your answer is not a fraction, it is a whole number

OpenStudy (anonymous):

4?

OpenStudy (misty1212):

no not 4? just 4

OpenStudy (anonymous):

-4?

OpenStudy (misty1212):

no \[24\div6=4\] you had it right the first time q

OpenStudy (anonymous):

ok so the awnser would be B

OpenStudy (anonymous):

???

OpenStudy (misty1212):

idk does B have a 4 as the first coordinate?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@misty1212 can you help me with one more